OpenStudy (anonymous):

approximately by what percentage does the volume of a sphere increase if it's surface area is increased 2%? See formulas below

OpenStudy (anonymous):

\[V = \frac{ 4 }{ 3 }\pi r^3\] \[S = 4\pi r^2\]

OpenStudy (anonymous):

answer is 3%, how do you get it though

OpenStudy (anonymous):

3% increase**

OpenStudy (anonymous):

what method are you estimating with? it really depends on the initial size of the sphere...

OpenStudy (anonymous):

are you learning differentials...?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{ dS }{ dt } = +0.02\]

OpenStudy (anonymous):

that is what we know

OpenStudy (anonymous):

radius is the problem